OpenStudy (anonymous):

Using the first three terms of the Maclaurin's series, what is the value of e^cos(π/2)?

OpenStudy (anonymous):

no, I dont write poems but I know how to write them

OpenStudy (anonymous):

The first three terms are as follows:\[f(x) = f(0) + f'(0)x + f''(0)x/2!\] in your case , f(x) = e^cos(pi/2) give it a try :)

OpenStudy (anonymous):

is it me or are you a bit tensed these days andrius?

OpenStudy (anonymous):

hope it's getting clear for you alma ^_^ give it a one more chance

OpenStudy (anonymous):

▓▒░╔ᴧᴨᴅᴙiᴜs╖░▒▓ it is my new signature you know, I am bussy making it better

OpenStudy (anonymous):

I know you're busy, but you've never been busy + tensed at the same time, and yes I have seen it, you've shown it to me and I like it, verfy mysterious.

OpenStudy (anonymous):

+ getting even more serious than ever.

OpenStudy (anonymous):

:) well, I guess I am growing up

OpenStudy (anonymous):

no andrius, it's not about growing up. Life's too precious to be taken seriously. You're serious at both the fun and serious moments which is not right. So many who are so young , yet so wise and so old than those who are the opposite.

OpenStudy (anonymous):

There's a difference of growing up and changing .

OpenStudy (anonymous):

between*

OpenStudy (anonymous):

my apologies alma, the third term is : \[(f''(0)/2)x^2\] instead of what I've written there.

OpenStudy (anonymous):

and here he goes running away again. Expected.

OpenStudy (anonymous):

nah, I am here ma buddy

OpenStudy (anonymous):

no you're not here, the andrius I know is vanishing already.

OpenStudy (anonymous):

relax, I am ok, the same I was yesterday, maybe more happy.... I am learning every day, I am smiling right now actually =D ^_^

OpenStudy (anonymous):

I am relaxed, always been and always have. Smiles sometimes tend to act as masks you know. From the tone of your words, you are not .

OpenStudy (anonymous):

ok, now I am relaxed and happy inside, calm inside, couldn't be better..

OpenStudy (anonymous):

why now?

OpenStudy (anonymous):

now as in this time, like today

OpenStudy (anonymous):

i sometimes express myself in a wrong way

OpenStudy (anonymous):

what was wrong with the others day then?

OpenStudy (anonymous):

I was confused

OpenStudy (anonymous):

you don't have to be. Damn, read your email.